### v3

# arXiv:2501.18323v3[math.SP] 02 Mar 2026

Abstract.
We study the approximation of eigenvalues for the Laplace-Beltrami operator on closed Riemannian manifolds in the class $\mathcal{M}$, characterized by bounded Ricci curvature, a lower bound on the injectivity radius, and an upper bound on the diameter. We use an $(\epsilon,\rho)$-approximation of the manifold by a weighted graph, as introduced by Burago et al. By adapting their methods, we prove that as the parameters $\epsilon,\rho$ and the ratio $\frac{\epsilon}{\rho}$ approach zero, the $k$-th eigenvalue of the graph Laplacian converges uniformly to the $k$-th eigenvalue of the manifold’s Laplacian for each $k$.
Key words and phrases:
Ricci curvature, spectrum of laplacian, graph laplacian
1991 Mathematics Subject Classification:
53C21,58J50, 58J60,35J05
1. Introduction
The approximation of eigenvalues and eigenfunctions of the Laplacian on manifolds using graphs and simplices is a fundamental problem in spectral geometry and has extensive applications in theoretical and applied fields. Dodziuk, Patodi, and Petrunin introduced frameworks connecting Riemannian structures with triangulations and polyhedral approximations of manifolds, providing an early foundation for discrete Laplacian analysis in [14], [15], [23]. Later works by Fujiwara [17] and Mantuano [21] on spectral convergence on manifolds under general geometric conditions have provided much of the conceptual basis for modern graph-based spectral approximation. A notable work by Belkin-Niyogi [3] established probabilistic convergence for random point samples.
In the deterministic setting, a major advance made by Burago, Ivanov, and Kurylev [8] established uniform spectral convergence of the graph Laplacian to the Laplace-Beltrami operator for closed Riemannian $n$-manifolds under uniform sectional curvature bounds, a lower injectivity radius bound, and an upper diameter bound. Aubry [2] obtained a similar result on approximating spectral data using isometrically immersed graphs in manifolds. Burago et al. [9] and Lu [20] also extended their framework to the $\rho$-Laplacian on metric measure spaces, including those with boundaries or certain singularities; further generalizations to vector bundles appear in [7]. Some recent related developments include [16], [10], [19], [12] and, [6]. Applications of the spectral convergence of graph Laplacians also arise in manifold learning and numerical computation, see for instance [[4],[18],[24],[25]].
The methods of Burago et al. [8] rely on estimates on the Jacobian of the exponential map derived from sectional curvature bounds. However, many fundamental results in Riemannian geometry, such as compactness theorems, eigenvalue bounds, and volume comparison, hold under the weaker assumption of Ricci curvature bounds. In this work, we establish a generalization of the spectral convergence theorem of [8] to the broader class $\mathcal{M}(n,\lambda,D,i_{0})$ of closed $n$-dimensional smooth Riemannian manifolds satisfying the following conditions:
$|Ric|\leq\lambda,\quad diam\leq D,\quad{\rm and}\quad inj\geq i_{0}$
where injinj and diamdiam denote the injectivity radius and diameter of the manifold, respectively. Anderson proved that $\mathcal{M}(n,\lambda,D,i_{0})$ is precompact in the $C^{1,\alpha}$-topology via harmonic coordinates in [1], thereby extending Cheeger’s finiteness theorem beyond its sectional curvature assumptions. The central contribution of our approach is to avoid dependencies on sectional curvature using estimates on the Jacobian of the exponential map derived from Anderson’s precompactness theorem. Thus, adapting the techniques in [8] of approximating manifolds using weighted graphs dependent on $\epsilon,\rho>0$, we show that as $\epsilon,\rho$ and the ratio $\frac{\epsilon}{\rho}$ tend to 0, the $k$-th eigenvalue of the graph Laplacian converges uniformly to the $k$-th eigenvalue of the Laplace-Beltrami operator for every $k$.
Definition 1 ($(\epsilon,\rho)$-approximation).
Given any $\rho>\epsilon>0$, a finite graph $\Gamma$ with the vertex set $F=\{x_{i}\}_{i=1}^{N}$ and edges $e_{ij}$} is called an $(\epsilon,\rho)$-approximation of $(M,g)$ if the following conditions hold.
  • $F\subset M$. There exists a partition of $M$ into measurable subsets $\{V_{i}\}_{i=1}^{N}$ such that $V_{i}\subset B_{\epsilon}(x_{i})$ and $M=\bigcup B_{\epsilon}(x_{i})$ where $B_{\epsilon}(x_{i})$ is the ball in $M$ centered at $x_{i}$ with radius $\epsilon$.
  • Two vertices $x_{i}$ and $x_{j}$ are connected by an edge $e_{ij}$ if the Riemannian distance between them is less than $\rho.$
  • Measure on $\Gamma$: Let $\mu_{i}=\text{vol}(V_{i}).$ $F$ can then be equipped with a discrete measure $\mu=\sum_{i=1}^{N}\mu_{i}\delta_{x_{i}}$ where $\delta_{x_{i}}$ denotes the Dirac measure at $x_{i}.$
  • Weights on edges: Let $\nu_{n}$ be the volume of the unit ball in the Euclidean $n$-space. To an edge $e_{ij}$, assign the weight
    $w_{ij}=\frac{2(n+2)}{\nu_{n}\rho^{n+2}}\mu_{i}\mu_{j}.$ (1.1)
For any $\textit{M}\in\mathcal{M}(n,\lambda,D,i_{0})$, there exists a finite set of points $F$ that is $\epsilon$-dense in M. We can then use a partition obtained by the Voronoi decomposition to define an $(\epsilon,\rho)$-approximation of $(M,g)$. The weighted graph is denoted by $\Gamma(F,\rho,\mu).$ A weighted graph Laplacian $\Delta_{\Gamma}$ on $L^{2}(\Gamma)$ is defined as follows:
$(\Delta_{\Gamma}u)(x_{i})$ $=\frac{1}{\mu_{i}}\sum_{x_{i}\sim x_{j}}w_{ij}(u(x_{i})-u(x_{j}))$
$=\frac{2(n+2)}{\nu_{n}\rho^{n+2}}\sum_{x_{i}\sim x_{j}}\mu_{j}(u(x_{i})-u(x_{j})).$ (1.2)
where $x_{i}\sim x_{j}$ is $x_{i}$ and $x_{j}$ are connected by an edge.
The motivation behind the choice of the normalization constant is given in Section 2.3 of [8]. We note that $\Delta_{\Gamma}$ is a non-negative self-adjoint operator on $L^{2}(\Gamma)$ with eigenvalues $0\leq\lambda_{1}(\Gamma)\leq\lambda_{2}(\Gamma)....\leq\lambda_{N}(\Gamma)$. The eigenvalues of $\Delta_{M}$ are denoted by $0\leq\lambda_{1}(M)\leq\lambda_{2}(M)\leq\lambda_{3}(M)\leq$.... The main goal of this paper is to prove the following theorem.
Theorem 1.1.
Consider a manifold $M\in\mathcal{M}(n,\lambda,D,i_{0})$ and an $(\epsilon,\rho)$-approximation $\Gamma$ of $M$ such that $\rho<i_{0}$. Let $\lambda_{k}(M)$, $\lambda_{k}(\Gamma)$ denote the $k$-th eigenvalue of $\Delta_{M}$ and $\Delta_{\Gamma}$ respectively. Then there exists a constant $C_{\mathcal{M}}>0$ such that for any $\rho,\frac{\epsilon}{\rho}<\frac{1}{C_{\mathcal{M}}}$,
$|\lambda_{k}(\Gamma)-\lambda_{k}(M)|\leq C_{\mathcal{M}}\left(\frac{\epsilon}{\rho}+\rho\right)\lambda_{k}(M)+\rho C_{\mathcal{M}}\lambda^{\frac{3}{2}}_{k}(M),\quad\forall\ k\leq|\Gamma|.$
In [11], Cheng proved that if the Ricci curvature of $M$ is bounded below by $\lambda$ and the diameter is bounded above by $D$, then the $k$-th eigenvalue of $\Delta_{M}$ is bounded above by a constant depending only on $k$, $n$, $\lambda$ and $D$. Hence, for any $M\in\mathcal{M}(n,\lambda,D,i_{0})$, there exists a uniform positive constant $\lambda_{\mathcal{M},k}$ such that $\lambda_{k}\leq\lambda_{\mathcal{M},k}.$ Consequently, we have the following corollary.
Corollary 1.2.
Consider a manifold $M\in\mathcal{M}(n,\lambda,D,i_{0})$ and an $(\epsilon,\rho)$-approximation $\Gamma$ of $M$ such that $\rho<i_{0}$. Let $\lambda_{k}(M)$, $\lambda_{k}(\Gamma)$ denote the $k$-th eigenvalue of $\Delta_{M}$ and $\Delta_{\Gamma}$ respectively. Then for each $k$, there exists a constant $C_{\mathcal{M},k}>0$ depending on $\mathcal{M}$ and $k$ such that for any $\rho,\frac{\epsilon}{\rho}<\frac{1}{C_{\mathcal{M},k}}$,
$|\lambda_{k}(\Gamma)-\lambda_{k}(M)|\leq C_{\mathcal{M},k}\left(\frac{\epsilon}{\rho}+\rho\right).$
Consequently, for every fixed $k$, $\lambda_{k}(\Gamma)$ converges uniformly to $\lambda_{k}(M)$ for all $M\in\mathcal{M}(n,\lambda,D,i_{0})$ as $\rho\to 0$ and $\frac{\epsilon}{\rho}\to 0.$
We now state the approximation of eigenfunctions of the Laplace-Beltrami operator by using the discretization scheme introduced above. We work with the discretization map $P$ (Definition 3) that connects functions on the manifold with functions on the discrete space and the interpolation map $I$ (Definition 7) which reconstructs a function on the manifold from discrete data.
Theorem 1.3.
Let $\lambda=\lambda_{j}(M)$ be an eigenvalue of $\Delta_{M}$ with multiplicity $m$, such that
$\lambda_{k-1}<\lambda_{k}=\lambda=\lambda_{k+m-1}<\lambda_{k+m}.$
Let $\delta_{\lambda}=\min\{1,\lambda_{k}-\lambda_{k-1},\lambda_{k+m}-\lambda_{k+m-1}\}$. Let $u_{k},\dots,u_{k+m-1}$ be orthonormal eigenvectors corresponding to $\lambda_{k}(\Gamma),\dots,\lambda_{k+m-1}(\Gamma)$.
Then there exist orthonormal eigenfunctions $g_{k},\dots,g_{k+m-1}$ of $\Delta_{M}$ corresponding to $\lambda$ such that for all $j=k,\dots,k+m-1$ and $\rho<\delta_{\lambda}C_{\mathcal{M},k}^{-1}$,
$\left\lVert u_{j}-Pg_{j}\right\rVert^{2}\leq C_{\mathcal{M},k}\delta_{\lambda}^{-2}\left(\frac{\epsilon}{\rho}+\rho\right)$ (1.3)
and
$\left\lVert g_{j}-Iu_{j}\right\rVert^{2}\leq C_{\mathcal{M},k}\delta_{\lambda}^{-2}\left(\frac{\epsilon}{\rho}+\rho\right).$ (1.4)
Idea of the proof:
By Rauch’s comparison theorem, if the absolute value of the sectional curvature is bounded by $K$, the Jacobian $J_{x}(v)$ of the exponential map at $x$ in the direction of $v$ satisfies
$\frac{1}{1+CnK|v|^{2}}\leq J_{x}(v)\leq 1+CnK|v|^{2},\quad\forall\ |v|<i_{0}.$
This inequality is a crucial component of the eigenvalue and eigenfunction approximation methods in [8]. However, this type of bound can not be easily obtained for bounded Ricci curvature. Although the authors in [8] noted the possibility of extending their methods to manifolds with Ricci curvature bounds, the generalization is not straightforward. We use Anderson’s precompactness theorem in [1] and derive appropriate estimates for the Jacobian in Lemma 2.2 that allow us to adopt the methods in [8] in the case of bounded Ricci curvature.
We prove Theorem 1.1 using the Min–Max principle. In Section 3, a discretization map from $L^{2}(M)$ to $L^{2}(\Gamma)$ is used to relate the discrete Dirichlet energy to the energy functional on $M$. Then by controlling the Rayleigh quotient on the image under the discretization map of the span of the first $k$ eigenfunctions of $\Delta_{M}$, these estimates yield an upper bound for $(\lambda_{k}(\Gamma)-\lambda_{k}(M))$.
The lower bound for $(\lambda_{k}(\Gamma)-\lambda_{k}(M))$ is obtained in Section 4. Using an interpolation map, we regularize elements of $L^{2}(\Gamma)$ into $C^{0,1}(M)$. This map is an approximate inverse of the discretization map. We then establish relations between the derivative of the interpolation map and the discrete Dirichlet energy. Considering the Rayleigh quotient restricted to the image of the interpolation map of the span of the first $k$ eigenfunctions of $\Delta_{\Gamma}$, and using our interpolation estimates, we obtain the desired lower bound.
2. Some inequalities and Average dispersion
Given a Riemannian manifold $(M,g)$ and a point $x\in M$, the exponential map at $x$ is denoted by $\exp_{x}$. Let $u$ be a unit vector in $T_{x}M.$ For $0<t<i_{0}$, using the normal polar coordinates $(u,t)$ in $T_{x}M$, we can write the volume element $\text{{vol}}_{g}$ as
$\exp_{x}^{*}\operatorname{vol}_{g}=J(u,t)t^{n-1}dtdu.$
where $J(u,t)$ is called the Jacobian determinant at $(u,t)$. In the sequel, we require suitable bounds $\lvert J(u,t)\rvert$ under the assumed geometric conditions. We first establish an estimate for the Laplacian of the distance function under an upper Ricci curvature bound.
Lemma 2.1 (Laplacian comparison of the distance function).
Let $(M,g)$ be a Riemannian n-manifold satisfying $\operatorname{Ric}\leq\lambda$, and $\operatorname{inj}(M)\geq i_{0}>0$. Fix $p\in M$ and let $r(x)=d(p,x)$ be the distance function from $p$. Then at $p\in B_{i_{0}/2}(x)\setminus\{x\}$,
$\Delta r(p)\geq\frac{n-1}{r(p)}-\lambda\,r(p).$
Proof.
Let $\gamma:[0,r(p)]\to M$ be the unique unit-speed minimizing geodesic from $x$ to $p$. One can write in polar coordinates
$g=dr^{2}+g_{r}.$
Let $H_{r}$ denote the Hessian of the distance function $r$. Then the fundamental equation (see [22] Chapter 2, section 5.2) can be written as
$\partial_{r}g=\partial_{r}g_{r}=2H_{r}$
and
$\partial_{r}H_{r}-H_{r}^{2}=-\operatorname{R}(.,\partial_{r},\partial_{r},.).$
By a taking trace and using the curvature assumption $\operatorname{Ric}\leq\lambda$, we have
$\partial_{r}\Delta r\geq\operatorname{tr}(H_{r}^{2})-\lambda\geq\frac{(\Delta r)^{2}}{n-1}-\lambda.$ (2.1)
The last inequality follows from the Cauchy-Schwarz inequality. Therefore,
$\partial_{r}\left(\Delta_{r}-\frac{n-1}{r}\right)$ $=\partial_{r}\Delta r+\frac{n-1}{r^{2}}-2\frac{\Delta r}{r}$
$\geq\frac{1}{n-1}(\Delta r)^{2}-\lambda+\frac{n-1}{r^{2}}-2\frac{\Delta r}{r}$ (2.2)
$=\frac{1}{n-1}\left(\Delta r-\frac{n-1}{r}\right)^{2}-\lambda$
$\geq-\lambda.$ (2.3)
Let $f(r)=\Delta r-\frac{n-1}{r}.$ Then,
$f(r)-f(\epsilon)$ $=\int_{\epsilon}^{r}f^{\prime}(s)ds$
$\geq-\int_{\epsilon}^{r}\lambda ds$
$=-\lambda(r-\epsilon).$ (2.4)
From [22] Chapter 5, section 7
$\lim_{r\to 0}\left(H_{r}-\frac{1}{r}g_{r}\right)=0.$
Taking the trace we have,
$\lim_{\epsilon\rightarrow 0}\left(\Delta r(\gamma(\epsilon))-\frac{n-1}{\epsilon}\right)=0.$
Hence,
$f(r)\geq-\lambda r.$
Which implies,
$\Delta r(p)\geq\frac{n-1}{r(p)}-\lambda r(p).$
Lemma 2.2.
There exists a constant $c_{\mathcal{M}}>0$ such that for any $M\in\mathcal{M}(n,\lambda,D,i_{0})$ and $x\in M$,
$|1-J(u,t)|\leq c_{\mathcal{M}}t,\quad\forall\ u\in S^{n-1},\quad\forall\ t\leq\frac{i_{0}}{2}.$
Proof.
Consider a polar coordinate decomposition on $B_{r}(x)$, the geodesic ball of radius $r\leq\frac{i_{0}}{2}$ centered at $x$ in $(M,g)\in\mathcal{M}.$ Then, from [22], Chapter 9, section 1 we have
$d\operatorname{vol}=\tilde{J}(u,t)du\wedge d\operatorname{vol}_{n-1}$
and
$\partial_{r}\tilde{J}=\tilde{J}\Delta r.$ (2.5)
where $\tilde{J}(u,t)=J(u,t)t^{n-1}.$ By differentiating (2.5), we obtain
$\Delta r=\frac{n-1}{r}+\frac{J^{\prime}(u,r)}{J(u,r)}.$ (2.6)
From Lemma 2.1 and the Laplace comparison theorem (see [22], Chapter 9, section 1.2) for $\operatorname{Ric}\geq-\lambda$, we have
$\frac{n-1}{t}-\lambda t\leq\Delta r(\gamma(t))\leq(n-1)\sqrt{\lambda}\coth(\sqrt{\lambda}t).$
Using (2.6),
$\lambda tJ(u,t)\leq J^{\prime}(u,t)\leq(n-1)\sqrt{\lambda}\coth(\sqrt{\lambda}t)J(u,t)-\frac{n-1}{t}J(u,t).$
From the Taylor series expansion of the $\coth$ function, we have $\lim_{t\rightarrow 0}\left(\sqrt{\lambda}\coth(\sqrt{\lambda}t)-\frac{1}{t}\right)=0$ and $J(u,t)\leq\left(\frac{\sinh(\sqrt{\lambda}t}{\sqrt{\lambda}t}\right)^{n-1}.$ Using $|J(u,t)|\leq C(n,\lambda,i_{0})$, we obtain a uniform upper bound $c_{\mathcal{M}(n,\lambda,D,i_{0})}$ on the derivative of $J(u,t)$ on $B_{x}(r)$ for $r\leq\frac{i_{0}}{2}.$
$|J^{\prime}(u,t)|\leq c_{\mathcal{M}},\quad\forall p=\exp_{x}(tu)\in B_{\frac{i_{0}}{2}}(x),\quad\forall M\in\mathcal{M}.$ (2.7)
From the proof of the Bishop-Gromov volume comparison theorem, we know that $J(u,0)=1$ for any $u\in S^{n-1}.$ Using mean value theorem and (2.7) we have
$|J(u,t)-1|=|J(u,t)-J(u,0)|\leq c_{\mathcal{M}}t.$
Definition 2.
Let $\textit{M}\in\mathcal{M}(n,\lambda,D,i_{0})$ and $B_{r}(x)$ be a ball in M centered at $x$ of radius $r$. Let $f:\textit{M}\to\mathbb{R}$ be a smooth function. The average dispersion $E_{r}(f)$ is defined as
$E_{r}(f)=\int_{\textit{M}}\int_{B_{r}(x)}|f(y)-f(x)|^{2}dydx.$
We recall that $H^{1}(M)$ is the space of all functions in $L^{2}(M)$ such that their first-order partial derivatives are also in $L^{2}(M)$. The following lemma gives an upper bound of $E_{r}(f)$.
Lemma 2.3.
For $M\in\mathcal{M}(n,\lambda,D,i_{0})$, let $0<r<\frac{i_{0}}{2},f\in H^{1}(\textit{M})$. Then
$E_{r}(f)\leq\frac{\nu_{n}r^{n+2}}{n+2}(1+c_{\mathcal{M}}r)\left\lVert df\right\rVert^{2}.$
Proof.
Since smooth functions are dense in $H^{1}(M)$, it is enough to prove the above inequality for smooth functions. For any $r<\frac{i_{0}}{2},$ the exponential map at x restricted to $B_{r}(x)$ is a diffeomorphism onto its image. Using polar co-ordinates $(u,t)$ on $T_{x}M$,
$\int_{B_{r}(x)}\lvert f(y)-f(x)\rvert^{2}dy=\int_{u\in S^{n-1}}\int_{t=0}^{r}\lvert f(exp_{x}(tu))-f(exp_{x}(0))\rvert^{2}J(u,t)t^{n-1}dtdu.$ (2.8)
Substituting the upper bound of $J(u,t)$ from Lemma 2.2 in equation (2.8),
$\int_{B_{r}(x)}|f(y)-f(x)|^{2}dy\leq(1+c_{\mathcal{M}}r)\int_{u\in S^{n-1}}\int_{t=0}^{r}\lvert f(exp_{x}(tu))-f(exp_{x}(0))\rvert^{2}t^{n-1}dtdu.$
Now,
$\int_{u\in S^{n-1}}\int_{t=0}^{r}\lvert f(exp_{x}(tu))-f(exp_{x}(0))\rvert^{2}t^{n-1}dtdu=\int_{B_{r}(0)\subset T_{x}\textit{M}}|f(exp_{x}(v))-f(x)|^{2}dv.$
By using the techniques for establishing an upper bound of the same integral in Lemma 3.3 ([8]), we have
$\int_{B_{r}(0)\subset T_{x}\textit{M}}|f(exp_{x}(v))-f(x)|^{2}dv\leq\frac{\nu_{n}r^{n+2}}{n+2}\left\lVert df\right\rVert^{2}_{L^{2}}.$
Combining this with the upper bound of the Jacobian determinant, the required inequality is obtained. ∎
In [5], Berger showed that there exists a constant $c_{n}>0$ depending only on $n$ such that
$\text{vol}(B_{r}(x))\geq c_{n}r^{n},\quad\forall r<i_{0}.$ (2.9)
Its values have been computed in [13]. We establish a Poincaré-type inequality using this estimate.
Lemma 2.4.
[…]
$\int_{V}|f(x)-a|^{2}dx\leq\frac{1}{c_{n}(r-\epsilon)^{n}}E_{r}(f,V).$
Proof.
The above inequality is obtained in Lemma 3.4 in [8] where an upper bound on the sectional curvature is used to obtain a lower bound on $\text{{vol}}(B_{r}(x))$. If $\text{{vol}}(B_{r}(x))\geq c_{n}r^{n}$, then the proof follows using the same techniques. ∎
3. An upper bound for $\lambda_{k}(\Gamma)$
We find an upper bound of the eigenvalues of $\Delta_{\Gamma}$ in terms of the eigenvalues of $\Delta_{M}$ for $M\in\mathcal{M}(n,\lambda,D,i_{0})$ using the inequalities proved in the previous section. Consider $M\in\mathcal{M}(n,\lambda,D,i_{0})$ and $\epsilon>0$ such that $0<\epsilon<\rho<\frac{i_{0}}{2}$. Let $\Gamma$ be an $(\epsilon,\rho)$-approximation of $(M,{\text{vol}})$ using a decomposition $\{V_{i}\}_{1}^{N}$. We define the discretization map that assigns to each vertex of the graph the average value of the function on $V_{i}$, the measurable set that contains the vertex.
Definition 3 (Discretization map).
Let $\mu_{i}={\text{vol}}(V_{i})$. Then $P:L^{2}(M)\to L^{2}(\Gamma)$ be a map defined as
$Pf(x_{i})=\frac{1}{\mu_{i}}\int_{V_{i}}fd\mu.$
The map $P^{*}:L^{2}(\Gamma)\to L^{2}(M)$ is defined as
$P^{*}(u)=\sum_{i=1}^{N}u(x_{i})\chi_{V_{i}}.$
We see some immediate properties of the discretization map in the following lemma.
Lemma 3.1.
Let $M\in\mathcal{M}(n,\lambda,D,i_{0}),f\in L^{2}(M)$ and $u\in L^{2}(\Gamma).$ Then
  1. (1)
    $P:L^{2}(M)\to L^{2}(\Gamma)$ is a bounded linear operator : $\left\lVert Pf\right\rVert\leq N\left\lVert f\right\rVert_{2}.$
  2. (2)
    $P^{*}$ is norm-preserving, i.e. $\left\lVert P^{*}(u)\right\rVert_{L^{2}(M)}=\left\lVert u\right\rVert_{L^{2}(\Gamma)}$.
  3. (3)
    $P^{*}$ is the adjoint of the discretization map $P$, i.e. $\langle f,P^{*}(u)\rangle_{L^{2}(M)}=\langle P(f),u\rangle_{L^{2}(\Gamma)}$.
Lemma 3.2.
If $0<\epsilon<\frac{\rho}{n}<\frac{i_{0}}{2n}$ and ${\text{vol}}(B_{r}(x))\geq c_{n}r^{n}$ for $\epsilon<r<\rho$ then
$\left\lVert f-P^{*}Pf\right\rVert^{2}_{L^{2}}\leq\frac{4n\nu_{n}}{c_{n}}(1+c_{\mathcal{M}}r)\epsilon^{2}\left\lVert df\right\rVert^{2}_{L^{2}}.$
Proof.
Let $V$ be a measurable subset of $M$ and $\chi_{V}$ denote the characteristic function on $V.$ According to definition,
$\left\lVert f-P^{*}Pf\right\rVert^{2}_{L^{2}}$ $=\int_{M}|f(x)-P^{*}Pf(x)|^{2}dx$
$=\int_{M}|f(x)-\sum_{i=1}^{N}Pf(x_{i})\chi_{V_{i}}(x)|^{2}dx$
$=\sum_{i=1}^{N}\int_{V_{i}}|f(x)-Pf(x_{i})|^{2}dx.$
From Lemma 2.4, for all $r$ such that $0<\epsilon<r<\rho$ and for all $1\leq i\leq N$,
$\int_{V_{i}}|f(x)-Pf(x_{i})|^{2}dx\leq\frac{1}{c_{n}(r-\epsilon)^{n}}E_{r}(f,V_{i}).$
From Lemma 2.3,
$\left\lVert f-P^{*}Pf\right\rVert^{2}_{L^{2}}\leq\frac{1}{c_{n}(r-\epsilon)^{n}}E_{r}(f)\leq\frac{\nu_{n}}{c_{n}(n+2)}\left(\frac{r}{r-\epsilon}\right)^{n}r^{2}(1+c_{\mathcal{M}}r)\left\lVert df\right\rVert^{2}_{L^{2}}.$
Putting $r=n\epsilon$, we have $\left(\frac{r}{r-\epsilon}\right)^{n}=\left(\frac{n}{n-1}\right)^{n}<4.$ Hence,
$\left\lVert f-P^{*}Pf\right\rVert^{2}_{L^{2}}\leq\frac{4n\nu_{n}}{c_{n}}(1+c_{\mathcal{M}}r)\epsilon^{2}\left\lVert df\right\rVert^{2}_{L^{2}}.$ (3.1)
As $P^{*}$ is norm-preserving, the next corollary follows immediately using the triangle inequality.
Corollary 3.3.
For any $f\in H^{1}(M),$
$|\left\lVert Pf\right\rVert-\left\lVert f\right\rVert_{L^{2}}|^{2}\leq\frac{4n\nu_{n}}{c_{n}}(1+c_{\mathcal{M}}r)\epsilon^{2}\left\lVert df\right\rVert^{2}_{L^{2}}.$
Proof.
$|\left\lVert Pf\right\rVert-\left\lVert f\right\rVert_{L^{2}}|^{2}\leq|\left\lVert P^{*}Pf\right\rVert-\left\lVert f\right\rVert|^{2}\leq\frac{4n\nu_{n}}{c_{n}}(1+c_{\mathcal{M}}r)\epsilon^{2}\left\lVert df\right\rVert^{2}_{L^{2}}.$
Next, we define the discrete Dirichlet energy and study its relation to the Dirichlet energy functional defined on $H^{1}(M).$
Definition 4 (Discrete Dirichlet energy).
Let $\Gamma=\Gamma(F,\rho,\mu)$ be an $(\epsilon,\rho)$- approximation of $M\in\mathcal{M}(n,\lambda,D,i_{0}).$ For $u\in L^{2}(\Gamma)$, the Dirichlet energy on $\Gamma$ is given by
$\left\lVert\delta(u)\right\rVert^{2}=\frac{n+2}{\nu_{n}\rho^{n+2}}\sum_{i}\sum_{j:x_{i}\sim x_{j}}\mu_{i}\mu_{j}|u(x_{j})-u(x_{i})|^{2}.$ (3.2)
Lemma 3.4.
Let $\Gamma$ be a $(\epsilon,\rho)$-approximation of $(M,{\text{vol}})$. For $f\in H^{1}(M)$,
$\left\lVert\delta(Pf)\right\rVert^{2}\leq(1+c_{\mathcal{M}}(\rho+2\epsilon))\left(1+\frac{2\epsilon}{\rho}\right)^{n+2}\left\lVert df\right\rVert^{2}_{L^{2}}.$
Proof.
From Lemma 4.3(2) in [8], we derive
$\left\lVert\delta(Pf)\right\rVert^{2}$ $\leq\frac{n+2}{\nu_{n}\rho^{n+2}}\sum_{i}\sum_{j:x_{j}\sim x_{i}}\int_{V_{i}}\int_{V_{j}}|f(y)-f(x)|^{2}dydx$
$=\frac{n+2}{\nu_{n}\rho^{n+2}}\int_{M}\int_{U(x)}|f(y)-f(x)|^{2}dydx.$ (3.3)
where $U(x)=\bigcup_{j:x_{j}\sim x_{i}}V_{j}$ in (3.3). We also observe that $U(x)\subseteq B_{\rho+2\epsilon}(x).$
Hence,
$\left\lVert\delta(Pf)\right\rVert^{2}\leq\frac{n+2}{\nu_{n}\rho^{n+2}}E_{\rho+2\epsilon}(f).$ (3.4)
Using Lemma 2.3 and from (3.4), we have
$\left\lVert\delta(Pf)\right\rVert^{2}$ $\leq(1+c_{\mathcal{M}}(\rho+2\epsilon))\frac{(\rho+2\epsilon)^{n+2}}{\rho^{n+2}}\left\lVert df\right\rVert^{2}_{L^{2}}$
$=(1+c_{\mathcal{M}}(\rho+2\epsilon))\left(1+\frac{2\epsilon}{\rho}\right)^{n+2}\left\lVert df\right\rVert^{2}_{L^{2}}.$
Now we can calculate an upper bound of the eigenvalue of the graph Laplacian operator using the lemmas proved in this section.
Theorem 3.5.
Let $0<\epsilon<\rho\leq\frac{i_{0}}{2}\text{, }M\in\mathcal{M}(n,\lambda,D,i_{0})$ and $\Gamma$ be a $(\epsilon,\rho)$-approximation of $(M,{\text{vol}})$. Let $\lambda_{k}(M)$ and $\lambda_{k}(\Gamma)$ denote the k-th eigenvalues of $\Delta_{M}$ and $\Delta_{\Gamma}$ respectively. There exist positive constants $C_{n}$ and $C_{\mathcal{M}}$ depending only on $n$ and $\mathcal{M}$ respectively such that for any $\rho$, $\frac{\epsilon}{\rho}<\frac{1}{C_{n}}$ ,
$\lambda_{k}(\Gamma)\leq\left(1+C_{n}\frac{\epsilon}{\rho}+C_{\mathcal{M}}\rho\right)\lambda_{k}(M)+\rho C_{\mathcal{M}}\lambda^{\frac{3}{2}}_{k}(M).$
Proof.
We use the min-max principle to prove the theorem. Hence it is sufficient to find a linear subspace $L\subset L^{2}(\Gamma)$ with dimL=kL=k such that the Rayleigh quotient restricted to $L$ given by $sup_{u\in L\setminus\{0\}}\frac{\left\lVert\delta u\right\rVert^{2}}{\left\lVert u\right\rVert^{2}}$ is less than or equal to the right-hand side of the above inequality.
Let $W\subset H^{1}(M)$ be the linear span of orthonormal eigenfunctions of $\Delta_{M}$ corresponding to eigenvalues $\lambda_{1}(M)\leq\lambda_{2}(M)\leq...\leq\lambda_{k}(M).$ For any $f\in W,$
$\left\lVert df\right\rVert^{2}_{L^{2}}=\langle-\Delta_{M}f,f\rangle\leq\lambda_{k}(M)\left\lVert f\right\rVert^{2}_{L^{2}}.$
From Corollary 3.3 and considering $\rho<\frac{1}{c_{\mathcal{M}}}$,
$\left\lVert Pf\right\rVert\geq\left\lVert f\right\rVert_{L^{2}}-2\epsilon\sqrt{\frac{n\nu_{n}}{c_{n}}(1+\rho c_{\mathcal{M}})}\left\lVert df\right\rVert_{L^{2}}\geq\left(1-3\epsilon\sqrt{\frac{n\nu_{n}}{c_{n}}\lambda_{k}(M)}\right)\left\lVert f\right\rVert_{L^{2}}.$ (3.5)
Let $\lambda_{\mathcal{M},k}$ be an upperbound of $\lambda_{k}(M)$ for any $M\in\mathcal{M}.$ Then for $\epsilon<\frac{\sqrt{c_{n}}}{3\sqrt{n\nu_{n}\lambda_{\mathcal{M},k}}}$, $P|_{W}$ is injective. Hence, the dimension of the subspace $L=P(W)$ is $k.$ We pick $u\in L\setminus\{0\}$ and let $f\in W$ be such that $u=Pf.$ Then from equation (3.5),
$\left\lVert u\right\rVert^{2}\geq\left(1-3\epsilon\sqrt{\frac{n\nu_{n}}{c_{n}}\lambda_{k}(M)}\right)\left\lVert f\right\rVert^{2}_{L^{2}}.$ (3.6)
By Lemma 3.4 for $\epsilon<\frac{\rho}{2}$,
$\left\lVert\delta(u)\right\rVert^{2}\leq(1+2\rho c_{\mathcal{M}})\left(1+\frac{2\epsilon}{\rho}\right)^{n+2}\lambda_{k}(M)\left\lVert f\right\rVert^{2}_{L^{2}}.$ (3.7)
From inequalities (3.6) and (3.7), we have
$\lambda_{k}(\Gamma)\leq\frac{\left\lVert\delta u\right\rVert^{2}}{\left\lVert u\right\rVert^{2}}$ $\leq\frac{(1+2\rho c_{\mathcal{M}})\left(1+\frac{2\epsilon}{\rho}\right)^{n+2}\lambda_{k}(M)}{\left(1-3\epsilon\sqrt{\frac{n\nu_{n}}{c_{n}}\lambda_{k}(M)}\right)}.$
Using the Taylor series expansion of terms $\left(1+\frac{2\epsilon}{\rho}\right)^{n+2}$ and $\left(1-3\epsilon\sqrt{\frac{n\nu_{n}}{c_{n}}\lambda_{k}(M)}\right)^{-1}$ we obtain a constant $C_{\mathcal{M}}$ depending only on $\mathcal{M}$ and a constant $C_{n}$ depending on $n$ such that for any $\epsilon,\frac{\epsilon}{\rho}<\frac{1}{C_{n}}$ ,
$\lambda_{k}(\Gamma)\leq\left(1+C_{n}\frac{\epsilon}{\rho}+C_{\mathcal{M}}\rho\right)\lambda_{k}(M)+\rho C_{\mathcal{M}}\lambda^{\frac{3}{2}}_{k}(M).$
Remark 3.6.
We observe that it is enough to assume a lower bound on Ricci curvature to obtain an upper bound on $J(u,t)$, which will give us a lower bound on $\lambda_{k}(M)$ using $\lambda_{k}(\Gamma)$. One can apply Bishop- Gromov volume comparison theorem and conclude that for Ricci curvature bounded below by $\lambda$ and $t<i_{0}$, $J(u,t)\leq J_{\lambda}(t)$ where
$J_{\lambda}(t)=\left(1+\sum_{j\in\mathbb{N}}\frac{|\lambda|^{j}t^{2j}}{(2j+1)!}\right)^{n-1}.$ (3.8)
We can then follow the calculations of Theorem 3 and obtain a lower bound of $\lambda_{k}(M).$
4. Smoothing Operator
In this section, we study the smoothing operator defined in [8] and its properties, which will help us to obtain a lower bound for the eigenvalues of $\Delta_{\Gamma}$ using eigenvalues of $\Delta_{M}$.
Definition 5.
Fixing a positive $r<\rho<\frac{i_{0}}{2}$, the following functions are defined:
  1. (1)
    Smoothing function. Let $\psi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ be defined as
[…]
  2. (2)
    Kernel. Let $k_{r}:M\times M\to\mathbb{R}_{+}$ be defined as
    $k_{r}(x,y)=r^{-n}\psi(r^{-1}d(x,y)).$
    where $d(x,y)$ denotes the Riemannian distance between $x$ and $y$ in $M.$
  3. (3)
    Associated integral operator. Let $\Lambda_{r}^{0}:L^{2}(M)\to C^{0,1}(M)$ be given by
    $\Lambda_{r}^{0}f(x)=\int_{M}f(y)k_{r}(x,y)dy.$
  4. (4)
    Let $\theta\in C^{0,1}(M)$ be defined as
    $\theta=\Lambda_{r}^{0}(\mathbf{1}_{M}).$
By using the polar coordinate in $\mathbb{R}^{n}$, the following lemma can be easily established.
Lemma 4.1.
Let $p\in\mathbb{R}^{n}$ and $x,y\in M$. Let $\psi,k_{r}$ be defined as above. Then,
  1. (1)
    $\int_{\mathbb{R}^{n}}\psi(|p|)dt=1$, and
  2. (2)
    $|k_{r}(x,y)|\leq\frac{n+2}{\nu_{n}r^{n}}.$
  3. (3)
    ${\rm grad}k_{r}(.,y)(x)=\frac{n+2}{\nu_{n}r^{n+2}}\exp^{-1}(y)$
We derive the following approximations for $\theta$ using the bounds of the Jacobian determinant from Lemma 2.2.
Lemma 4.2.
For $0<r<\frac{i_{0}}{2}\text{ and }M\in\mathcal{M}(n,\lambda,D,i_{0})$, let $x\in M\text{ and }B_{r}(x)$ be a ball of radius r centered at x. Then $\theta(x)$ has the following bounds
$1-c_{\mathcal{M}}r\leq\theta(x)\leq 1+c_{\mathcal{M}}r.$
The next lemma gives a bound on the $L^{2}$ norm of the gradient of $\theta$.
Lemma 4.3.
Let $0<r<\rho<\frac{i_{0}}{2}\text{ and }M\in\mathcal{M}(n,\lambda,D,i_{0})$. Let $\nabla\theta$ denote the gradient of $\theta$. Then $|\nabla\theta(x)|\leq c_{\mathcal{M}}.$
Proof.
Let $A:S^{n-1}\to S^{n-1}$ be the antipodal map defined as A(u)=u.A(u)=-u. We note that the Jacobian matrix corresponding to $A$ is I-I, where $I$ is the identity matrix. By changing co-ordinates using $A$, we can show that
$\int_{u\in S^{n-1}}udu=0.$ (4.1)
One can compute the gradient of $\theta$ in polar coordinates as follows:
$\nabla\theta(x)=\frac{n+2}{\nu_{n}r^{n+2}}\int_{t=0}^{r}\int_{u\in S^{n-1}}tuJ(u,t)t^{n-1}dudt.$
From equation (4.1),
$\int_{u\in S^{n-1}}\int_{t=0}^{r}t^{n}ududt=\int_{u\in S^{n-1}}udu\int_{t=0}^{r}t^{n}dt=0.$
Hence,
$\int_{u\in S^{n-1}}\int_{t=0}^{r}tJ(u,t)t^{n-1}ududt=\int_{u\in S^{n-1}}\int_{t=0}^{r}t^{n}(J(u,t)-1)ududt$
From Lemma 2.2, we have
$|J(u,t)-1|$ $\leq tc_{\mathcal{M}}.$ (4.2)
Let $w\in T_{x}M$, $|w|=1$ be such that $\max\{\langle\triangledown\theta(x),v\rangle|v\in T_{x}M$, |v|=1}|v|=1\} is attained at $w$. One can then compute the gradient of $\theta$ in polar coordinates as follows:
$\lvert\nabla\theta(x)\rvert=\langle\nabla\theta,w\rangle$ $=\frac{n+2}{\nu_{n}r^{n+2}}\int_{u\in S^{n-1}}\int_{t=0}^{r}t^{n}\langle u,w\rangle\left(J(u,t)-1\right)dtdu$
$\leq\frac{n+2}{\nu_{n}r^{n+2}}\int_{u\in S^{n-1}}\int_{t=0}^{r}|\langle u,w\rangle||1-J(u,t)|t^{n}dtdu$
$\leq\frac{n+2}{\nu_{n}r^{n+2}}\int_{u\in S^{n-1}}\int_{t=0}^{r}c_{\mathcal{M}}t^{n+1}dtdu$
$\leq c_{\mathcal{M}}.$
Using the associated integral operator and $\theta$, we define the following function.
Definition 6.
Let $\Lambda_{r}:L^{2}(M)\to C^{0,1}(M)$ be defined as $\Lambda_{r}f=\theta^{-1}\Lambda_{r}^{0}f$.
We observe that $\Lambda_{r}f=f$, for all constant functions $f.$ In fact, $\Lambda_{r}f$ is bounded when $r$ is less than the injectivity radius as we see in the following lemma.
Lemma 4.4.
Let $0<r<\frac{i_{0}}{2}\text{ and }M\in\mathcal{M}(n,\lambda,D,i_{0})$. For all $f\in L^{2}(M)$ and $r<\frac{1}{c_{\mathcal{M}}}$, we have
$\left\lVert\Lambda_{r}f\right\rVert_{L^{2}}^{2}\leq\left(\frac{1+c_{\mathcal{M}}r}{1-c_{\mathcal{M}}r}\right)\left\lVert f\right\rVert_{L^{2}}^{2}.$
Proof.
Let $x\in M$. Then
$|\Lambda_{r}^{0}f(x)|^{2}$ $=|\int_{M}f(y)k_{r}(x,y)dy|^{2}$
$\leq\left(\int_{M}(\sqrt{k_{r}(x,y)})^{2}dy\right)\left(\int_{M}(f(y)\sqrt{k_{r}(x,y)})^{2}dy\right)$ (4.3)
$=\theta(x)\int_{M}|f(y)|^{2}k_{r}(x,y)dy.$
where (4.3) follows from Cauchy-Schwarz inequality.
So,
$|\Lambda_{r}f(x)|^{2}$ $=\theta(x)^{-2}|\Lambda_{r}^{0}f(x)|^{2}$
$\leq\theta(x)^{-1}\int_{M}|f(y)|^{2}k_{r}(x,y)dy$
$\leq(1-c_{\mathcal{M}}r)^{-1}\int_{M}|f(y)|^{2}k_{r}(x,y)dy.$
Integrating with respect to the volume measure on $M$,
$\left\lVert\Lambda_{r}f\right\rVert^{2}_{L^{2}}=\int_{M}|\Lambda_{r}f(x)|^{2}dx$ $\leq(1-c_{\mathcal{M}}r)^{-1}\int_{M}|f(y)|^{2}\int_{M}k_{r}(x,y)dydx$
$\leq(1-c_{\mathcal{M}}r)^{-1}(1+c_{\mathcal{M}}r)\left\lVert f\right\rVert^{2}_{L^{2}}$ (4.4)
where (4.4) follows from Lemma 4.2. ∎
When $r$ is less than half of the injectivity radius, the following lemma gives an upper bound using average dispersion on how far $\Lambda_{r}$ takes a function in $L^{2}(M).$
Lemma 4.5.
Let $0<r<\frac{i_{0}}{2}\text{ and }M\in\mathcal{M}(n,\lambda,D,i_{0})$. For all $f\in L^{2}(M)$ and $r<\frac{1}{c_{\mathcal{M}}}$,
$\left\lVert\Lambda_{r}f-f\right\rVert^{2}_{L^{2}}\leq\frac{n+2}{\nu_{n}r^{n}(1-c_{\mathcal{M}}r)}E_{r}(f).$
Proof.
We fix a coset representative $f:M\to\mathbb{R}$ of an element of $L^{2}(M).$
Let $a=f(x).$ Now,
$\Lambda_{r}f(x)-f(x)$ $=\Lambda_{r}f(x)-a$
$=\Lambda_{r}(f-a.\mathbf{1}_{M})(x)$
$=\theta^{-1}(x)\int_{B_{r}(x)}(f(y)-a)k_{r}(x,y)dy$
$=\theta^{-1}(x)\int_{B_{r}(x)}(f(y)-f(x))k_{r}(x,y)dy.$
By Cauchy-Schwarz inequality,
$|\Lambda_{r}f(x)-f(x)|^{2}$ $\leq\theta^{-2}(x)\left(\int_{B_{r}(x)}k_{r}(x,y)dy\right)\left(\int_{B_{r}(x)}|f(y)-f(x)|^{2}k_{r}(x,y)dy\right)$
$=\theta^{-1}(x)\int_{B_{r}(x)}|f(y)-f(x)|^{2}k_{r}(x,y)dy$
$\leq\frac{n+2}{\nu_{n}r^{n}(1-c_{\mathcal{M}}r)}\int_{B_{r}(x)}|f(y)-f(x)|^{2}dy$ (4.5)
where (4.5) follows from Lemma 4.1. Finally, by integrating on both sides, we get the result. ∎
In the following lemma, we obtain an upper bound for the norm of $d\Lambda_{r}(f)$ using the average dispersion $E_{r}f$ under Ricci curvature lower bound.
Lemma 4.6.
Let $0<r<\rho<\frac{i_{0}}{2}\text{ and }M\in\mathcal{M}(n,\lambda,D,i_{0})$. For every $f\in L^{2}(M),$
$\left\lVert d(\Lambda_{r}f)\right\rVert^{2}\leq\frac{n+2}{\nu_{n}r^{n+2}}\frac{(1+2c_{\mathcal{M}}r)^{2}}{(1-c_{\mathcal{M}}r)^{3}}E_{r}f.$
Proof.
Fixing a coset representative of $L^{2}(M)$ and denoting it by $f$, let $\tilde{f}=\Lambda_{r}f.$ For any $a\in\mathbb{R},$
$\tilde{f}(x)=a+\theta^{-1}(x)\int_{B_{r}(x)}(f(y)-a)k_{r}(x,y)dy.$
Using chain rule and differentiating $\tilde{f}$ we get,
$d_{x}\tilde{f}=\theta^{-1}(x)\int_{B_{r}(x)}(f(y)-a)d_{x}k_{r}(.,y)dy+d_{x}(\theta^{-1})\int_{B_{r}(x)}(f(y)-a)k_{r}(x,y)dy.$
Putting $a=f(x),$ we have
$d_{x}\tilde{f}=\theta(x)^{-1}A_{1}(x)+A_{2}(x).$ (4.6)
Here,
$A_{1}(x)=\int_{B_{r}(x)}(f(y)-f(x))d_{x}k_{r}(.,y)dy$
and
$A_{2}(x)=d_{x}(\theta^{-1})\int_{B_{r}(x)}(f(y)-f(x))k_{r}(x,y)dy.$
From Lemma 4.2, and using triangle inequality,
$\left\lVert d\tilde{f}\right\rVert_{L^{2}}\leq(1-c_{\mathcal{M}}r)^{-1}\left\lVert A_{1}\right\rVert_{L^{2}}+\left\lVert A_{2}\right\rVert_{L^{2}}.$
We evaluate each integral $A_{1}$ and $A_{2}$ separately to get our desired inequality.
Using Cauchy-Schwarz inequality,
$|A_{2}(x)|^{2}$ $\leq|d_{x}(\theta^{-1})|^{2}\left(\int_{B_{r}(x)}k_{r}(x,y)dy\right)\left(\int_{B_{r}(x)}|f(y)-f(x)|^{2}k_{r}(x,y)dy\right)$
$\leq|d_{x}(\theta^{-1})|^{2}\theta(x)\int_{B_{r}(x)}|f(y)-f(x)|^{2}k_{r}(x,y)dy$
$\leq\frac{n+2}{\nu_{n}r^{n}}\frac{|d_{x}\theta|^{2}}{\theta^{3}}\int_{B_{r}(x)}|f(y)-f(x)|^{2}dy$
$\leq\frac{(n+2)}{\nu_{n}r^{n}}\frac{c_{\mathcal{M}}^{2}}{(1-c_{\mathcal{M}}r)^{3}}\int_{B_{r}(x)}|f(y)-f(x)|^{2}dy.$
Integrating on $M$,
$\left\lVert A_{2}\right\rVert_{L^{2}}\leq\frac{c_{\mathcal{M}}r}{(1-c_{\mathcal{M}}r)^{\frac{3}{2}}}\sqrt{\frac{n+2}{\nu_{n}r^{n+2}}E_{r}(f)}.$ (4.7)
On the other hand, fixing $x\in M,$ let $w\in T_{x}M$ be such that $\text{{max}}\{\lvert\langle A_{1}(x),v\rangle\rvert:v\in T_{x}M,|v|=1\}$ is attained at $w.$
So,
$\lvert A_{1}(x)\rvert$ $=\langle A_{1}(x),w\rangle$
$=\frac{n+2}{2\nu_{n}r^{n+2}}\int_{B_{r}(x)}\left(f(y)-f(x)\right)\langle\exp^{-1}(y),w\rangle dy$
$=\frac{n+2}{2\nu_{n}r^{n+2}}\int_{t=0}^{r}\int_{u\in S^{n-1}}\left(f(\exp_{x}(tu))-f(\exp_{x}(0))\right)\langle tu,w\rangle t^{n-1}dudt.$
By Cauchy-Schwarz inequality,
$\lvert A_{1}(x)\rvert^{2}\leq\left(\frac{n+2}{2\nu_{n}r^{n+2}}\right)^{2}\left(\int_{B_{r}(0)}\lvert f(\exp_{x}(tu)-f(\exp_{x}(0))\rvert^{2}J(u,t)^{2}t^{n-1}dudt\right)\\ \left(\int_{B_{r}(0)}\langle tu,w\rangle^{2}t^{n-1}dudt\right).$
Now, $\frac{n+2}{\nu_{n}r^{n+2}}\int_{B_{r}(0)}\langle tu,w\rangle^{2}t^{n-1}dudt=1.$ Hence,
$\lvert A_{1}(x)\rvert^{2}$ $\leq\frac{n+2}{\nu_{n}r^{n+2}}\int_{B_{r}(0)}\lvert f(\exp_{x}(tu))-f(\exp_{x}(0))\rvert^{2}J(u,t)^{2}t^{n-1}dudt$
$\leq\frac{n+2}{\nu_{n}r^{n+2}}(1+c_{\mathcal{M}}r)\int_{B_{r}(0)}\lvert f(\exp_{x}(tu))-f(\exp_{x}(0))\rvert J(u,t)t^{n-1}dudt$
$=\frac{n+2}{\nu_{n}r^{n+2}}(1+c_{\mathcal{M}}r)\int_{B_{r}(x)}\lvert f(y)-f(x)\rvert^{2}dy.$
Integrating on $M$,
$\left\lVert A_{1}\right\rVert_{L^{2}}\leq\sqrt{\frac{n+2}{\nu_{n}r^{n+2}}(1+c_{\mathcal{M}}r)E_{r}f}.$ (4.8)
From (4.6), (4.7) and (4.8),
$\left\lVert df\right\rVert_{L^{2}}$ $\leq\sqrt{\frac{n+2}{\nu_{n}r^{n+2}}}\left(\frac{1+c_{\mathcal{M}}r}{1-c_{\mathcal{M}}r}+\frac{c_{\mathcal{M}}r}{(1-c_{\mathcal{M}}r)^{\frac{3}{2}}}\right)\sqrt{E_{r}f}$
$\leq\sqrt{\frac{n+2}{\nu_{n}r^{n+2}}}\frac{1+2c_{\mathcal{M}}r}{(1-c_{\mathcal{M}}r)^{\frac{3}{2}}}\sqrt{E_{r}f}.$
Putting the value of $f_{\lambda}(r)$ in the above inequality, the lemma follows. ∎
5. A lower bound for $\lambda_{k}(\Gamma)$
In this section, we compute an appropriate lower bound for the eigenvalues of $\Delta_{\Gamma}$ using the eigenvalues of $\Delta_{M}$ for the manifolds in $\mathcal{M}(n,\lambda,D,i_{0}).$
Definition 7 (Interpolation map).
Let $M\in\mathcal{M}(n,\lambda,D,i_{0})$ and $F$ be a finite $\epsilon$-net in $M$ such that $(F,\mu)\text{ }\epsilon$- approximates $(M,vol).$
Define $I:L^{2}(\Gamma)\to C^{0,1}(M)$ as
$Iu=\Lambda_{\rho-2\epsilon}P^{*}u.$
From Lemma 4.4 and the fact that $P^{*}$ preserves the norm, we have
$\left\lVert Iu\right\rVert_{L^{2}}\leq\left(\frac{1+c_{\mathcal{M}}(\rho-2\epsilon)}{1-c_{\mathcal{M}}(\rho-2\epsilon)}\right)\left\lVert u\right\rVert_{L^{2}}.$
Lemma 5.1.
Let $0<\epsilon<\rho<i_{0}$, $c_{\mathcal{M}}\rho<\frac{1}{4}$ and $\frac{2\epsilon}{\rho}<\frac{1}{n}.$ Consider $M\in\mathcal{M}(n,\lambda,D,i_{0})$ and $\Gamma$ be a $(\epsilon,\rho)$-approximation of $M$. Then
  1. (1)
    $|\left\lVert Iu\right\rVert_{L^{2}}-\left\lVert u\right\rVert|^{2}\leq\frac{3\rho^{2}}{1-c_{\mathcal{M}}\rho}\left\lVert\delta u\right\rVert^{2}.$
  2. (2)
    $\left\lVert dIu\right\rVert_{L^{2}}\leq\left(1-\frac{2\epsilon}{\rho}\right)^{-\frac{n}{2}-1}\frac{1+2c_{\mathcal{M}}\rho}{(1-c_{\mathcal{M}}\rho)^{\frac{3}{2}}}\left\lVert\delta u\right\rVert.$
Proof.
(1) Let $u\in L^{2}(\Gamma).$ Using norm preserving property of $P^{*}$ and triangle inequality,
$\lvert\left\lVert Iu\right\rVert_{L^{2}}-\left\lVert u\right\rVert\rvert\leq\left\lVert Iu-P^{*}u\right\rVert_{L^{2}}.$
Using Lemma 4.5, we have
$\left\lVert Iu-P^{*}u\right\rVert_{L^{2}}^{2}=\left\lVert\Lambda_{\rho-2\epsilon}P^{*}u-P^{*}u\right\rVert^{2}_{L^{2}}\leq\frac{n+2}{\nu_{n}(1-c_{\mathcal{M}}(\rho-2\epsilon))(\rho-2\epsilon)^{n}}E_{\rho-2\epsilon}(P^{*}u).$ (5.1)
On the other hand,
$\left\lVert\delta u\right\rVert^{2}=\frac{n+2}{\nu_{n}\rho^{n+2}}\int_{M}\int_{U(x)}\lvert P^{*}u(y)-P^{*}u(x)\rvert^{2}dydx,$
where $U(x)=\bigcup_{j:x_{j}\sim x_{i}}V_{j}$ such that $x\in V_{i}$. Since $U(x)\supset B_{\rho-2\epsilon}(x)$,
$\left\lVert\delta u\right\rVert^{2}$ $\geq\frac{n+2}{\nu_{n}\rho^{n+2}}\int_{M}\int_{B_{\rho-2\epsilon}(x)}\lvert P^{*}u(y)-P^{*}u(x)\rvert^{2}dydx$
$=\frac{n+2}{\nu_{n}\rho^{n+2}}E_{\rho-2\epsilon}(P^{*}u).$ (5.2)
Thus, from (5.1) and (5),
$\left\lVert Iu-P^{*}u\right\rVert_{L^{2}}^{2}\leq\frac{\rho^{n+2}}{(1-c_{\mathcal{M}}(\rho-2\epsilon))(\rho-2\epsilon)^{n}}\left\lVert\delta u\right\rVert^{2}\leq\frac{3\rho^{2}}{1-c_{\mathcal{M}}\rho}\left\lVert\delta u\right\rVert^{2}$ (5.3)
as $\frac{\rho^{n+2}}{(\rho-2\epsilon)^{n}}<3\rho^{2}$ for $\frac{2\epsilon}{\rho}<\frac{1}{n}$.
(2) Proceeding with direct computation,
$\left\lVert dIu\right\rVert^{2}_{L^{2}}$ $=\left\lVert d(\Lambda_{\rho-2\epsilon}P^{*}u)\right\rVert^{2}_{L^{2}}$
$\leq\frac{n+2}{\nu_{n}(\rho-2\epsilon)^{n+2}}\frac{(1+2c_{\mathcal{M}}\rho)^{2}}{(1-c_{\mathcal{M}}\rho)^{3}}E_{\rho-2\epsilon}(P^{*}u)$ (5.4)
$\leq\left(\frac{\rho}{\rho-2\epsilon}\right)^{n+2}\frac{(1+2c_{\mathcal{M}}\rho)^{2}}{(1-c_{\mathcal{M}}\rho)^{3}}\left\lVert\delta u\right\rVert^{2}$ (5.5)
where (5.4) follows from Lemma 4.6 and (5.5) follows from (5). ∎
Using the above lemma and the results obtained on the interpolation map, we are now ready to compute, under certain constraints, a lower bound for the eigenvalue of the graph Laplacian operator using the eigenvalue of the Laplace-Beltrami operator for all $M\in\mathcal{M}(n,\lambda,D,i_{0})$. Since $\lambda_{\mathcal{M},k}$ is an upper bound for all $k$-th eigenvalues without loss of generality, we consider $\lambda_{\mathcal{M},k}>3(1+4c_{\mathcal{M}}).$
Theorem 5.2.
For $0<\epsilon<\rho<\frac{i_{0}}{2}$, let $M\in\mathcal{M}(n,\lambda,D,i_{0}),$ and $\Gamma=(F,\mu)$ be an $(\epsilon,\rho)$- approximation of $(M,{\text{vol}}).$ Let $\lambda_{k}(M)$ and $\lambda_{k}(\Gamma)$ denote the k-th eigenvalues of $\Delta_{M}$ and $\Delta_{\Gamma}$ respectively. There exist positive constants $C_{n}$ and $C_{\mathcal{M}}$ depending solely on $n$ and $\mathcal{M}$, respectively, such that
$\lambda_{k}(\Gamma)\geq\left(1-C_{n}\frac{\epsilon}{\rho}-C_{\mathcal{M}}\rho\right)\lambda_{k}(M)-C_{\mathcal{M}}\rho\lambda_{k}^{\frac{3}{2}}(M).$
Proof.
If $\lambda_{k}(\Gamma)>\lambda_{k}(M)$ then the theorem follows immediately. So, we assume that $\lambda_{k}(\Gamma)\leq\lambda_{k}(M)$. We use the min-max principle again to prove the above inequality. Therefore, it is enough to show there is a subspace $L\subset H^{1}(M)$ with dimL=kL=k such that for any non-zero $f\in L$, $\frac{\|df\|_{L^{2}}^{2}}{\|f\|_{L^{2}}^{2}}$ is less than or equal to the right hand side of the above inequality. We consider the $k$-dimensional subspace $W$ of $L^{2}(\Gamma)$ which is the linear span of the first $k$ orthonormal eigenfunctions of $\Delta_{\Gamma}$. Let $u\in W$. Then we have $\left\lVert\delta u\right\rVert^{2}\leq\lambda_{k}(\Gamma)\left\lVert u\right\rVert^{2}.$ Using Lemma 5.1(1),
$\left\lVert Iu\right\rVert_{L^{2}}\geq\left(1-\rho\sqrt{\frac{3\lambda_{k}(\Gamma)}{1-c_{\mathcal{M}}\rho}}\right)\left\lVert u\right\rVert.$
Let $f=Iu.$ As $\lambda_{k}(\Gamma)\leq\lambda_{k}(M)$ we have,
$\left\lVert f\right\rVert^{2}_{L^{2}}\geq\left(1-\rho\sqrt{\frac{3\lambda_{k}(M)}{1-c_{\mathcal{M}}\rho}}\right)^{2}\left\lVert u\right\rVert^{2}$ (5.6)
From Lemma 5.1(2),
$\left\lVert df\right\rVert^{2}_{L^{2}}$ $\leq\left(1-\frac{2\epsilon}{\rho}\right)^{-n-2}\frac{(1+2c_{\mathcal{M}}\rho)^{2}}{(1-c_{\mathcal{M}}\rho)^{3}}\left\lVert\delta u\right\rVert^{2}$
$\leq\left(1-\frac{2\epsilon}{\rho}\right)^{-n-2}\frac{(1+2c_{\mathcal{M}}\rho)^{2}}{(1-c_{\mathcal{M}}\rho)^{3}}\lambda_{k}(\Gamma)\left\lVert u\right\rVert^{2}.$ (5.7)
From equations (5.6) and (5), we have
$\lambda_{k}(M)\leq\frac{\left\lVert df\right\rVert^{2}}{\left\lVert f\right\rVert^{2}}$ $\leq\left(1-\frac{2\epsilon}{\rho}\right)^{-n-2}\frac{(1+2c_{\mathcal{M}}\rho)^{2}}{(1-c_{\mathcal{M}}\rho)^{2}(1-c_{\mathcal{M}}\rho-4\rho\sqrt{(1-\rho c_{\mathcal{M}})\lambda_{k}(M)})}\lambda_{k}(\Gamma).$
For $\rho<\frac{1}{4c_{\mathcal{M}}}$ and $\frac{2\epsilon}{\rho}<1$ we have,
$\lambda_{k}(\Gamma)$ $\geq(1-\frac{2\epsilon}{\rho})^{n+2}(1-c_{\mathcal{M}}\rho)^{2}(1-c_{\mathcal{M}}\rho-4\rho\sqrt{(1-\rho c_{\mathcal{M}})\lambda_{k}(M)})(1+2c_{\mathcal{M}}\rho)^{-2}\lambda_{k}(M)$ (5.8)
$\geq\left(1-2(n+2)\frac{\epsilon}{\rho}\right)(1-2c_{\mathcal{M}}\rho)\left(1-c_{\mathcal{M}}\rho-4\rho\sqrt{\lambda_{k}(M)}\right)(1+5c_{\mathcal{M}}\rho)^{-1}\lambda_{k}(M)$
Now the theorem follows from the Taylor series expansion of $(1+5c_{\mathcal{M}}\rho)^{-1}.$
Proof of Theorem 1.1 :
The proof follows from Theorem 3 and Theorem 5.2.
6. Approximation of eigenfunctions
In this section, we investigate the approximation of eigenfunctions of $M\in\mathcal{M}(n,\lambda,D,i_{0})$ using similar arguments as in [8]. The following lemma shows that the discretization map and the interpolation map are almost inverses of each other.
Lemma 6.1.
  1. (1)
    Let $f\in H^{1}(M)$. Then for $\rho<\frac{1}{2c_{\mathcal{M}}}$,
    $\left\lVert IPf-f\right\rVert\leq C_{n}\rho\left\lVert df\right\rVert.$
  2. (2)
    Let $u\in L^{2}(\Gamma)$, $\rho\leq\frac{1}{3c_{\mathcal{M}}}$ and $\frac{\epsilon}{\rho}<\frac{1}{2n}$. Then,
    $\left\lVert PIu-u\right\rVert\leq C_{n}\rho\left\lVert\delta u\right\rVert$
where $C_{n}$ is a positive constant depending only on the dimension $n.$
Proof.
  1. (1)
    By straightforward computation,
    $\left\lVert IPf-f\right\rVert$ $=\left\lVert\Lambda_{\rho-2\epsilon}P^{*}Pf-f\right\rVert$
    $\leq\left\lVert\Lambda_{\rho-2\epsilon}(P^{*}Pf-f)\right\rVert+\left\lVert\Lambda_{\rho-2\epsilon}f-f\right\rVert.$ (6.1)
    We aim to find upper bounds of both the terms obtained above using the triangle inequality. From Lemma 4.4 and Lemma 3.2,
    $\left\lVert\Lambda_{\rho-2\epsilon}(P^{*}Pf-f)\right\rVert^{2}\leq\frac{4n\nu_{n}}{c_{n}}\frac{(1+c_{\mathcal{M}}\rho)^{2}}{1-c_{\mathcal{M}}\rho}\epsilon^{2}\left\lVert df\right\rVert_{L^{2}}^{2}.$ (6.2)
    Again from Lemma 4.5 and Lemma 2.3,
    $\left\lVert\Lambda_{\rho-2\epsilon}f-f\right\rVert^{2}\leq\frac{n+2}{\nu_{n}(1-c_{\mathcal{M}}\rho)(\rho-2\epsilon)^{n}}E_{\rho-2\epsilon}(f)\leq\frac{1+c_{\mathcal{M}}\rho}{1-c_{\mathcal{M}}\rho}\rho^{2}\left\lVert df\right\rVert^{2}.$ (6.3)
    From equations (6.1), (6.2) and (6.3), we have,
    $\left\lVert IPf-f\right\rVert\leq\frac{(1+c_{\mathcal{M}}\rho)}{\sqrt{1-c_{\mathcal{M}}\rho}}\left(1+\sqrt{\frac{n\nu_{n}}{c_{n}}}\right)\rho\left\lVert df\right\rVert.$
    As $c_{\mathcal{M}}\rho<\frac{1}{2}$, the first part of the lemma follows.
  2. (2)
    Since $P^{*}$ preserves norm, using the triangle inequality
    $\left\lVert PIu-u\right\rVert=\left\lVert P^{*}(PIu-u)\right\rVert\leq\left\lVert P^{*}PIu-Iu\right\rVert+\left\lVert Iu-P^{*}u\right\rVert.$ (6.4)
    Using Lemma 3.2, for $\rho<\frac{1}{c_{\mathcal{M}}}$,
    $\left\lVert P^{*}PIu-Iu\right\rVert^{2}\leq\frac{4n\nu_{n}}{c_{n}}(1+c_{\mathcal{M}}\rho)\epsilon^{2}\left\lVert dIu\right\rVert^{2}.$
    By Lemma 5.1 we have,
    $\left\lVert dIu\right\rVert^{2}_{L^{2}}\leq(1-\frac{2\epsilon}{\rho})^{-n-2}\frac{(1+2c_{\mathcal{M}}\rho)^{2}}{(1-c_{\mathcal{M}}\rho)^{3}}\left\lVert\delta u\right\rVert^{2}$
    Therefore, for $\rho<\frac{1}{c_{\mathcal{M}}}$ and $\frac{2\epsilon}{\rho}<\frac{1}{n}$,
    $\left\lVert P^{*}PIu-Iu\right\rVert\leq 2\sqrt{\frac{3n\nu_{n}}{c_{n}}}\frac{(1+2c_{\mathcal{M}}\rho)^{\frac{3}{2}}}{(1-c_{\mathcal{M}}\rho)^{\frac{3}{2}}}\epsilon\left\lVert\delta u\right\rVert.$ (6.5)
    From equation (5.3) we have,
    $\left\lVert Iu-P^{*}u\right\rVert\leq\frac{3\rho}{1-2c_{\mathcal{M}}\rho}\left\lVert\delta u\right\rVert.$ (6.6)
    Hence from (6.6) and (6.5),
    $\left\lVert PIu-u\right\rVert\leq 3\left(1+\sqrt{\frac{\nu_{n}}{c_{n}}}\right)\frac{(1+2c_{\mathcal{M}}\rho)^{\frac{3}{2}}}{(1-2c_{\mathcal{M}}\rho)^{\frac{3}{2}}}\rho\left\lVert\delta u\right\rVert.$
The second part of the lemma follows from the fact that $c_{\mathcal{M}}\rho<\frac{1}{3}.$
Let $I\subset\mathbb{R}$ be an interval, and define $H_{I}(M)$(respectively $H_{I}(X)$) as the subspace of $H^{1}(M)$ (respectively $H^{1}(X)$) spanned by eigenfunctions corresponding to eigenvalues in $I.$ We denote $H_{(-\infty,\lambda)}(M)$ by $H_{\lambda}(M).$ Let $\mathbb{P}_{I}:L^{2}(M)\to H_{I}(M)$ be the orthogonal projection. Analogous notation is used for projections from $L^{2}(\Gamma)$ to $H_{J}(X)$ and we write $\mathbb{P}_{\lambda}$ as shorthand for $\mathbb{P}_{(-\infty,\lambda)}$.
Lemma 6.2.
  1. (1)
    Let $\lambda>0$ and $f\in H_{\lambda}(M).$ Then
    $\left\lVert\delta(Pf)\right\rVert\geq\left(1-\sigma\right)\left\lVert df\right\rVert$
    where $\sigma=C_{\mathcal{M}}\rho+C_{n}\sqrt{\lambda}\rho+C_{n}\frac{\epsilon}{\rho}.$
  2. (2)
    Let $\lambda>0$ and $u\in H_{\lambda}(X).$ Then
    $\left\lVert d(Iu)\right\rVert\geq\left(1-\sigma\right)\left\lVert\delta u\right\rVert$
    where $\sigma=C_{\mathcal{M}}\rho+C_{n}\sqrt{\lambda}\rho+C_{n}\frac{\epsilon}{\rho}.$
Proof.
  1. (1)
    As $\mathbb{P}_{\lambda}$ is non-increasing with respect to the Dirichlet energy,
    $\left\lVert d(IPf)\right\rVert_{L^{2}}\geq\left\lVert d(\mathbb{P}_{\lambda}IPf)\right\rVert_{L^{2}}\geq\left\lVert df\right\rVert_{L^{2}}-\left\lVert d(\mathbb{P}_{\lambda}IPf-f)\right\rVert_{L^{2}}.$
    Given that $f\in H_{\lambda}(M)$, Lemma 6.1 implies for $\rho<\frac{1}{3c_{\mathcal{M}}}$,
    $\left\lVert d(\mathbb{P}_{\lambda}IPf-f)\right\rVert_{L^{2}}\leq\sqrt{\lambda}\|\mathbb{P}_{\lambda}(IPf-f)\|$ $\leq C_{n}\sqrt{\lambda}\rho\left\lVert df\right\rVert.$
    Hence,
    $\left\lVert d(IPf)\right\rVert_{L^{2}}\geq\left(1-C_{n}\sqrt{\lambda}\rho\right)\left\lVert df\right\rVert.$ (6.7)
    For $\rho<\frac{1}{4c_{\mathcal{M}}}$ and $\frac{\epsilon}{\rho}<\frac{1}{2n}$, from Lemma 5.1 there exists a constant $C_{n},C_{\mathcal{M}}>0$ such that
    $\left\lVert d(IPf)\right\rVert\leq(1+C_{n}\frac{\epsilon}{\rho}+C_{\mathcal{M}}\rho)\left\lVert\delta Pf\right\rVert.$ (6.8)
    Thus, from (6.7) and (6.8),
    $\left\lVert\delta(Pf)\right\rVert$ $\geq\left(1-\sigma\right)\left\lVert df\right\rVert$
    where $\sigma=C_{\mathcal{M}}\rho+C_{n}\sqrt{\lambda}\rho+C_{n}\frac{\epsilon}{\rho}.$
  2. (2)
    Applying the techniques from the previous part and Lemma 6.1, for $\rho<\frac{1}{3c_{\mathcal{M}}}$ we obtain
    $\left\lVert\delta(PIu)\right\rVert_{L^{2}}\geq\left(1-C_{n}\sqrt{\lambda}\rho\right)\left\lVert\delta u\right\rVert.$ (6.9)
    From Lemma 3.4, for $\frac{\epsilon}{\rho}<\frac{1}{C_{n}}$
    $\left\lVert\delta(PIu)\right\rVert\leq(1+C_{n}\frac{\epsilon}{\rho}+C_{\mathcal{M}}\rho)\left\lVert d(Iu)\right\rVert.$ (6.10)
    From (6.9) and (6.10),
    $\left\lVert d(Iu)\right\rVert$ $\geq\frac{1-C_{n}\sqrt{\lambda}\rho}{(1+C_{n}\frac{\epsilon}{\rho}+C_{\mathcal{M}}\rho)}\left\lVert\delta u\right\rVert$
    $\geq(1-\sigma)\left\lVert\delta u\right\rVert$
    where $\sigma=C_{\mathcal{M}}\rho+C_{n}\sqrt{\lambda}\rho+C_{n}\frac{\epsilon}{\rho}.$
Let $\{f_{k}\}_{1}^{\infty}$ and $\{u_{k}\}_{1}^{N}$ be orthonormal eigenvectors of $\Delta_{M}$ and $\Delta_{X}$ respectively. The next lemma approximates eigenfunctions of $\Delta_{M}$, which has been proved using Lemma 7.3 from [8] with analogous inequalities and approximations in our setting.
Lemma 6.3.
Let $0<\epsilon<\rho<\frac{i_{0}}{2}$, $k\in\mathbb{N}$ and $a\in\mathbb{R}_{+}$
  1. (1)
    Let $\lambda=\lambda_{k}(M).$ Then
    $\left\lVert Pf_{k}-\mathbb{P}_{\lambda_{k+a}}Pf_{k}\right\rVert^{2}\leq\frac{1}{a}C_{\mathcal{M},k}\left(\frac{\epsilon}{\rho}+\rho\right).$
    Also
    $\left\lVert\delta(Pf_{k}-\mathbb{P}_{\lambda+a}Pf_{k})\right\rVert^{2}\leq C_{\mathcal{M},k}\left(1+\frac{1}{a}\right)\left(\frac{\epsilon}{\rho}+\rho\right).$
  2. (2)
    Let $\lambda=\lambda_{k}(X).$ Then
    $\left\lVert Iu_{k}-\mathbb{P}_{\lambda+a}Iu_{k}\right\rVert^{2}\leq\frac{1}{a}C_{\mathcal{M},k}\left(\frac{\epsilon}{\rho}+\rho\right).$
    Also
    $\left\lVert d(Iu_{k}-\mathbb{P}_{\lambda+a}Iu_{k}\right\rVert^{2}\leq C_{\mathcal{M},k}\left(1+\frac{1}{a}\right)\left(\frac{\epsilon}{\rho}+\rho\right).$
Proof.
  1. (1)
    From the proof of Theorem 3.5, we have dimLL= $k$ if $\epsilon<\frac{1}{C_{\mathcal{M},k}}$. Let $Q$ denote the discrete Dirichlet energy form on $L^{2}(\Gamma)$ and $\{\lambda_{j}^{L}:1\leq j\leq k\}$ be the eigenvalues of $Q$ restricted to $L.$
    For $f\in W$, from Corollary 3.3
    $\left(1-C_{n}\left(\frac{\epsilon}{\rho}+c_{\mathcal{M}}\rho\right)\right)\left\lVert f\right\rVert\leq\left\lVert Pf\right\rVert\leq\left(1+C_{n}\left(\frac{\epsilon}{\rho}+c_{\mathcal{M}}\rho\right)\right)\left\lVert f\right\rVert$ (6.11)
    and
    $\left\lVert\delta(Pf)\right\rVert\leq C_{n}\left(\frac{\epsilon}{\rho}+c_{\mathcal{M}}\rho\right)\left\lVert df\right\rVert.$
    By minimax principle, for $\rho<\frac{1}{C_{n}c_{\mathcal{M}}}$,
    $\lambda_{j}^{L}\leq\lambda_{j}(M)+\lambda_{\mathcal{M},k}\left(1+C_{n}\left(\frac{\epsilon}{\rho}+c_{\mathcal{M}}\rho\right)\right)$ (6.12)
    Let $Q^{\prime}:L^{2}(\Gamma)\rightarrow\mathbb{R}$ be defined as
    $Q^{\prime}(u)=Q(\mathbb{P}_{\lambda+a}(u))+\lambda\left\lVert u-\mathbb{P}_{\lambda+a}(u)\right\rVert^{2}.$
    We observe that $Q^{\prime}$ is a quadratic form and $Q^{\prime}\leq Q.$ Hence, for every $j\leq m$ and $V\subset L^{2}(\Gamma)$ such that dimV=jV=j,
    $\sup_{v\in V\setminus\{0\}}\frac{Q^{\prime}(v)}{\left\lVert v\right\rVert^{2}}\geq\min\{\lambda,\lambda_{j}(\Gamma)\}.$
    Let $\{\lambda_{j}^{\prime}:1\leq j\leq k\}$ be the eigenvalues of $Q^{\prime}|_{L}.$ Using the above equality and the minimax principle $\lambda^{\prime}_{j}\geq\min\{\lambda,\lambda_{j}(\Gamma)\}$ for all $j\leq k.$
    By Theorem 1.1,
    $\lambda^{\prime}_{j}\geq\lambda_{j}(M)-C_{\mathcal{M},k}\left(\frac{\epsilon}{\rho}+\rho\right).$
    From (6.12),
    $\lambda_{j}^{L}-\lambda_{j}^{\prime}\leq C_{\mathcal{M},k}\left(\frac{\epsilon}{\rho}+\rho\right)$
    for all $j\leq k.$ Hence, for every $u\in L$,
    $Q(u)-Q^{\prime}(u)\leq C_{\mathcal{M},k}\left(\frac{\epsilon}{\rho}+\rho\right).$ (6.13)
    Let $u^{\prime}=u-\mathbb{P}_{\lambda+a}u.$ Since $Q(u^{\prime})\geq(\lambda+a)\left\lVert u^{\prime}\right\rVert^{2},$
    $Q(u)-Q^{\prime}(u)=Q(u)-Q^{\prime}(u^{\prime})\geq\frac{a}{\lambda+a}Q(u^{\prime}).$ (6.14)
    From (6.13) and (6.14),
    $Q(u^{\prime})\leq\frac{\lambda+a}{a}C_{\mathcal{M},k}\left(\frac{\epsilon}{\rho}+\rho\right)\left\lVert u\right\rVert^{2}.$
    Hence,
    $\left\lVert u^{\prime}\right\rVert^{2}\leq\frac{1}{a}C_{\mathcal{M},k}\left(\frac{\epsilon}{\rho}+\rho\right)\left\lVert u\right\rVert^{2}.$
    Substituting $u=Pf_{k}$, and observing from (6.11) that for sufficiently small $\epsilon,\rho,\frac{\epsilon}{\rho}$, $\left\lVert Pf_{k}\right\rVert<2$, we obtain the required expression.
  2. (2)
    The second assertion follows analogously using Lemma 5.1 and appropriate assumptions for $c_{\mathcal{M}}$ and $\lambda_{\mathcal{M},k}$ as above.
The following is a generalization of Lemma 7.4 from [8].
Lemma 6.4.
Let $\epsilon<\mathcal{C}_{\mathcal{M},k}^{-1}$ and $\mathcal{C}_{\mathcal{M},k}>1$.
  1. (1)
    Let $\lambda=\lambda_{k}(M)$ such that $0<\alpha\leq\beta\leq\gamma\leq 1$ and there are no eigenvalues of $\Delta_{\Gamma}$ in $(\lambda+\alpha,\lambda+\beta).$ Then
    $\left\lVert Pf_{k}-\mathbb{P}_{(\lambda-\gamma,\lambda+\alpha]}Pf_{k}\right\rVert^{2}\leq C_{\mathcal{M},k}\alpha\gamma^{-1}+C_{\mathcal{M},k}\beta^{-1}\gamma^{-1}\left(\frac{\epsilon}{\rho}+\rho\right).$
  2. (2)
    Let $\lambda=\lambda_{k}(\Gamma)$ and $\alpha,\beta,\gamma>0$ such that $\alpha\leq\beta\leq\gamma\leq 1$ and there are no eigenvalues of $\Delta_{M}$ in $(\lambda+\alpha,\lambda+\beta).$ Then
    $\left\lVert Iu_{k}-\mathbb{P}_{(\lambda-\gamma,\lambda+\alpha]}Iu_{k}\right\rVert^{2}\leq C_{\mathcal{M},k}\alpha\gamma^{-1}+C_{\mathcal{M},k}\beta^{-1}\gamma^{-1}\left(\frac{\epsilon}{\rho}+\rho\right).$
Proof.
  1. (1)
    Let $Q$ denote the discrete Dirichlet energy on $L^{2}(\Gamma)$ and $u=Pf_{k}$. We can express $u$ as a decomposition of orthogonal vectors
    $u=u_{0}+u_{-}+u_{+}$
    where $u_{0}\in H_{(\lambda-\gamma,\lambda+\alpha)}(X)$, $u_{-}\in H_{(-\infty,\lambda-\gamma]}(X)$ and $u_{+}\in H_{[\lambda+\alpha,\infty)}(X)$. From Lemma 6.3,
    $\left\lVert u_{+}\right\rVert^{2}\leq\frac{1}{\beta}C_{\mathcal{M},k}\left(\frac{\epsilon}{\rho}+\rho\right)$ (6.15)
    and the Dirichlet energy for is bounded above by
    $Q(u_{+})\leq\frac{1}{\beta}C_{\mathcal{M},k}\left(\frac{\epsilon}{\rho}+\rho\right).$
    Using appropriate bounds on $c_{\mathcal{M}}$ and $\lambda_{\mathcal{M},k}$ from Lemma 6.2,
    $Q(u)\geq\left(1-C_{\mathcal{M},k}\left(\frac{\epsilon}{\rho}+\rho\right)\right)\lambda.$
    Hence,
    $Q(u_{0})+Q(u_{-})=Q(u)-Q(u_{+})\geq\lambda-\frac{1}{\beta}C_{\mathcal{M},k}\left(\frac{\epsilon}{\rho}+\rho\right).$
    By the min-max principle,
    $Q(u_{0})\leq(\lambda+\alpha)\left\lVert u_{0}\right\rVert^{2}$
    and
    $Q(u_{-})\leq(\lambda-\gamma)\left\lVert u_{-}\right\rVert^{2},$
    Thus,
    $\lambda-\frac{1}{\beta}C_{\mathcal{M},k}\left(\frac{\epsilon}{\rho}+\rho\right)\leq\lambda(\left\lVert u_{0}\right\rVert^{2}+\left\lVert u_{-}\right\rVert^{2})+\alpha\left\lVert u_{0}\right\rVert^{2}\gamma\left\lVert u_{-}\right\rVert^{2}.$
    Also, from (6.11),
    $\left\lVert u_{0}\right\rVert^{2}\leq\left\lVert Pf_{k}\right\rVert^{2}\leq 1+C_{\mathcal{M},k}\left(\frac{\epsilon}{\rho}+\rho\right).$
    Hence,
    $\lambda-\frac{1}{\beta}C_{\mathcal{M},k}\left(\frac{\epsilon}{\rho}+\rho\right)\leq\lambda(1+\sigma)+\alpha(1+\sigma)-\gamma\left\lVert u_{-}\right\rVert^{2}$
    where $\sigma=C_{\mathcal{M},k}\left(\frac{\epsilon}{\rho}+\rho\right)$. Hence for $\rho<\frac{\alpha\gamma}{C_{\mathcal{M},k}}$,
    $\left\lVert u_{-}\right\rVert^{2}\leq\beta^{-1}\gamma^{-1}C_{\mathcal{M},k}\left(\frac{\epsilon}{\rho}+\rho\right)+\alpha\gamma^{-1}(1+C_{\mathcal{M},k})$
    Using (6.15), we derive the required assertion.
Following similar methods as in the first part and taking appropriate bounds for $c_{\mathcal{M}}$ and $\lambda_{\mathcal{M},k}$, the second assertion can be obtained. ∎
Theorem 6.5.
  1. (1)
    Let $\lambda=\lambda_{k}(M)$ and let $f_{k}$ be corresponding unit-norm eigenfunction of $\Delta_{M}.$ Then for every $\gamma\in(0,1)$,
    $\left\lVert Pf_{k}-\mathbb{P}_{(\lambda-\gamma,\lambda+\gamma)}Pf_{k}\right\rVert^{2}\leq C_{\mathcal{M},k}\gamma^{-2}\left(\frac{\epsilon}{\rho}+\rho\right).$
  2. (2)
    Let $\lambda=\lambda_{k}(M)$ and let $f_{k}$ be corresponding unit-norm eigenfunction of $\Delta_{M}.$ Then for every $\gamma\in(0,1)$,
    $\left\lVert Iu_{k}-\mathbb{P}_{(\lambda-\gamma,\lambda+\gamma)}Iu_{k}\right\rVert^{2}\leq C_{\mathcal{M},k}\gamma^{-2}\left(\frac{\epsilon}{\rho}+\rho\right).$
Proof.
Substituting $\alpha=\beta=C_{\mathcal{M},k}\left(\frac{\epsilon}{\rho}+\rho\right)^{\frac{1}{2}}\gamma$ in Lemma 6.4 and using the fact that $\gamma^{-2}\geq 1$ we get the required expression. ∎
Theorem 6.6.
Let $\lambda=\lambda_{j}(M)$ be an eigenvalue of $\Delta_{M}$ with multiplicity $m$, such that
$\lambda_{k-1}<\lambda_{k}=\lambda=\lambda_{k+m-1}<\lambda_{k+m}.$
Let $\delta_{\lambda}=\min\{1,\lambda_{k}-\lambda_{k-1},\lambda_{k+m}-\lambda_{k+m-1}\}$. Let $u_{k},\dots,u_{k+m-1}$ be orthonormal eigenvectors corresponding to $\lambda_{k}(\Gamma),\dots,\lambda_{k+m-1}(\Gamma)$.
Then there exist orthonormal eigenfunctions $g_{k},\dots,g_{k+m-1}$ of $\Delta_{M}$ corresponding to $\lambda$ such that for all $j=k,\dots,k+m-1$ and $\rho<\delta_{\lambda}C_{\mathcal{M},k}^{-1}$,
$\left\lVert u_{j}-Pg_{j}\right\rVert^{2}\leq C_{\mathcal{M},k}\delta_{\lambda}^{-2}\left(\frac{\epsilon}{\rho}+\rho\right)$ (6.16)
and
$\left\lVert g_{j}-Iu_{j}\right\rVert^{2}\leq C_{\mathcal{M},k}\delta_{\lambda}^{-2}\left(\frac{\epsilon}{\rho}+\rho\right).$ (6.17)
Proof.
The proof of the above theorem is analogous to that of Theorem 4 in [8], utilizing the results of Lemma 6.4 and Lemma 6.2. ∎